At what temperature will toluene have a vapor pressure of 517 torr. The normal boiling point of toluene is 110.6C at 1 atm pressure with the molar enthalpy of vaporization being 35.2 kJ/mol
Answer #1

T1 = 110.6 oC
=(110.6 + 273)K
= 383.6 K
P1 = 7.6*10^2 torr [This is atmospheric pressure. This is the pressure at normal boiling point]
P2 = 5.17*10^2 torr
Hvap = 35.2 KJ/mol
= 35200.0 J/mol

use:
ln(P2/P1) = (Hvap/R)*(1/T1 - 1/T2)
ln(517.0/7.6*10^2) = (35200.0/8.314)*(1/383.6 - 1/T2)
-0.385276 = 4233.822468*(1/383.6 - 1/T2)
T2 = 371
= (371-273) oC
= 98 oC

Answer: 98 oC